# When Is Degree Enough? Bounds on Degree-Eigenvector Misalignment in Assortative Structured Networks

### v1

@@ 6 Heavy-tailed degrees and eigenvector localization
For assortativity, the affine constants in our moment bounds are pinned down by endpoint-degree dispersion. Newman assortativity is a normalized covariance of degrees across edges, and $\phi_{p,q}$ simply correlates the $p$-degree at the tail of an edge with the $q$-degree at its head (out/in, in/in, etc.). A degree-preserving swap therefore cannot change the endpoint-degree multisets. It can only change ‘who is paired with whom’ across edges, i.e. the cross-edge covariance itself. Which is exactly what assortativity records.
The normalization $\nu_{p,q}=1/(\sigma_{T}\sigma_{H})$ turns the raw degree-product covariance into a dimensionless correlation, where $\sigma_{T}$ and $\sigma_{H}$ are the standard deviations of the tail- and head-side endpoint-degree samples. Intuitively, this fixes the ‘units’ of mixing: when endpoint degrees are highly dispersed, the same absolute change in the degree-product sum produces a smaller standardized change in $\phi_{p,q}$. Along our rewiring trajectory, $\sigma_{T}$ and $\sigma_{H}$ (hence $\nu_{p,q}$) remain invariant, so the scale is set once and for all by the degree distribution, while rewiring acts only through changes in the cross-edge covariance.
Finally, because each swap re-pairs only a constant number of endpoints, the per-swap movement in $\phi_{p,q}$ is driven by the upper tail: the largest increments come from reassigning edges incident to rare high-degree vertices, which generate the biggest changes in the degree-product sum. Put together, this is why assortativity is usually slow to move: each swap re-pairs only a few endpoints, and the statistic averages over all edges, so a visible change typically requires many coordinated re-pairings. In heavy-tailed networks this picture becomes more uneven. High-degree endpoints are rare, so most swaps touch only moderate degrees and produce tiny changes. But when a swap does involve a rare high-degree node, the degree–product sum can jump, and so can $\phi_{p,q}$. In particular, heavier tails tend to worsen the worst-case affine constants in our moment bounds: even though greater dispersion reduces the standardized effect of a typical swap (via $\nu_{p,q}=1/(\sigma_{T}\sigma_{H})$), the presence of more extreme high-degree endpoints increases the size of the occasional jumps that drive the upper envelope. Therefore, in the assortativity case, the affine constants that bound the first two moments are controlled by endpoint-degree dispersion (through $\sigma_{T}\sigma_{H}$ and $\nu_{p,q}$) together with how much mass sits in the upper tail, since these two ingredients determine both the typical increment size and the occasional large jumps by which degree-preserving swaps can move $\phi_{p,q}$.
Communities.
For community structure, the affine constants are shaped by the chosen partition because $\phi_{\mathrm{com}}$ is a mass-transfer statistic, i.e. it measures how much edge mass lies within blocks relative to across blocks. A degree-preserving swap cannot change how many stubs each vertex contributes, so the only way to increase $\phi_{\mathrm{com}}$ is to re-route existing cross-block edges into within-block positions while keeping every vertex’s degree fixed. In that sense the rewiring is a constrained transport problem: one is trying to ‘pour’ as much mass as possible into the diagonal blocks of the adjacency matrix without changing the row/column sums.
How fast this transfer can proceed depends on two interacting ingredients. First, the partition geometry—in particular the number of communities and the resulting block sizes—controls the combinatorial supply of admissible within-block endpoints. With many communities (hence smaller blocks), within-block room is scarcer, so the process saturates sooner. With fewer, larger (or more balanced) blocks, there are simply more eligible within-block pairs to absorb diverted mass. Second, the degree tail matters because high-degree vertices dominate the supply and demand of edge mass: most of what can be moved is carried by the few large-degree nodes. If these hubs are distributed across blocks in a way that gives them abundant within-block partners, cross-block mass can be converted quickly. If hub degree mass is concentrated in blocks with limited within-block capacity, gains in $\phi_{\mathrm{com}}$ stall early even if many swaps remain available.
This is why, in the community case, the bounding affine constants depend on the interaction between the number of communities (which sets within-block capacity) and the tail of the degree distribution (which concentrates the movable mass): the former determines how much within-block ‘space’ exists, and the latter determines how much mass is trying to flow into it.
Core-periphery.
Core-periphery is similar in spirit to communities in that we relocate edge mass under degree preservation, but the target is asymmetric. The statistic $\phi_{\mathrm{cp}}$ rises mainly by suppressing periphery–periphery adjacency: we drain the $\mathcal{L}\!\to\!\mathcal{L}$ block and redirect that mass so that periphery endpoints connect through the core $\mathcal{H}$. Since out- and in-stub counts are fixed vertexwise, rewiring cannot change how much periphery stub mass exists. It can only change where it lands. Increasing $\phi_{\mathrm{cp}}$ is therefore a progressive transfer of $\mathcal{L}\!\to\!\mathcal{L}$ edges into $\mathcal{L}\!\to\!\mathcal{H}$, $\mathcal{H}\!\to\!\mathcal{L}$, and/or $\mathcal{H}\!\to\!\mathcal{H}$ locations.
The affine bounding constants are governed by the split $(\mathcal{H},\mathcal{L})$ and by how long this drainage can continue before capacity constraints bite. Under degree preservation, the core has a fixed aggregate ability to absorb redirected periphery edge-ends. If that capacity is insufficient, a nontrivial residue of $\mathcal{L}\!\to\!\mathcal{L}$ mass is unavoidable. Degree dispersion matters here too: in heavy-tailed networks a few very high-degree core vertices act as ‘large-capacity hubs’, creating more room to reattach periphery edge-ends to $\mathcal{H}$ before saturation.
The core-periphery statistic can have high leverage: a modest number of well-chosen swaps can rapidly funnel walk mass through $\mathcal{H}$, a structural asymmetry to which the Perron vector is particularly responsive. Which is why the affine bounds in the core-periphery case depend on the interaction between $(\mathcal{H},\mathcal{L})$ split and the heaviness of the degree tail, since the two together determine how much periphery mass can be redirected through the core and how sharply paths can be concentrated therein.
Cycles.
@@
For $k$-cycles, $\phi_{k\text{-}\mathrm{cyc}}$ is motif-based and the relevant rewiring is inherently local: a single accepted swap can create a new short cycle (or extend an incipient one) by rearranging only a constant number of edges, without any global reorganization of mixing patterns. This locality is precisely why cycles are a useful stress-test for eigenvector robustness: inserting a small motif in a small region can induce a noticeable, localized eigenvector distortion even when coarse global summaries move slowly. Degree dispersion still matters at this combinatorial level: high-degree vertices offer a large menu of neighbors, so when a swap touches a hub it is often easier to ‘close’ a cycle around it simply because there are many more admissible ways to complete the required adjacency pattern under degree preservation.
Longer cycles are typically costlier because a $k$-cycle requires coordinating adjacency among $O(k)$ vertices. In practice one often has to stage the motif over multiple accepted swaps, so the per-swap increase in $\phi_{k\text{-}\mathrm{cyc}}$ tends to shrink as $k$ grows. Heavy tails, however, reshape this cost: hubs supply abundant ‘attachment points’, so long cycles can be stitched together with far less coordination than in thin-tailed graphs. In this sense, the longer the target cycle, the more the tail thickness matters as hubs are precisely what make many kinds of long-cycle constructions feasible.
6.1 Heavy tails and the moment bounds
A unifying theme across all four statistics is that the heaviness of the degree tail governs how much structural change one can generate with a fixed degree-preserving perturbation budget. When the degree distribution is heavy-tailed, a small set of hubs carries a disproportionate share of edge-ends. This concentration creates two generic effects that recur throughout our constructions. First, it generates high leverage: swaps that involve hubs can move the statistic substantially, because much of the relevant mass transfer, edge pairing, or motif closure is mediated by these few vertices. Second, it makes the evolution lumpy: most swaps that touch moderate-degree nodes do little, but the occasional hub-involving swap can produce a visible jump. In worst-case terms, this is precisely what inflates the affine constants—heavier tails expand the range of one-step (and short-horizon) changes achievable under the same degree-preserving operator-norm budget, and hence widen the envelope of possible eigenvector rotations.
This lens helps reconcile the four cases. For assortativity, heavy tails amplify the rare but large covariance jumps created by re-pairing edges incident to hubs (even as tail dispersion also affects the normalization). For communities and core–periphery, heavy tails place much of the ‘movable mass’ in a few vertices, so the speed of mass transfer is governed by whether the partition/core split gives those hubs enough admissible destinations before capacity saturates. For cycles, tail heaviness matters combinatorially: hubs offer many local routes, making it easier not only to close short cycles but also to assemble and stitch together longer cycles, since hubs supply abundant attachment points for coordinating adjacency across many vertices.
There is, however, a further (and subtler) channel through which tail heaviness can enter the moment bounds: it may also affect the spectral sensitivity parameters themselves. Hub-centered rewiring typically builds multiple forms of structure around the same small set of vertices (with many low-degree neighbors acting as ‘feeders’ or attachment points), and this can make distinct eigenmodes less separable. In the language of the Stewart–Sun bound, this corresponds to potential increases in the distortion factor $\kappa$ (a more ill-conditioned eigenbasis) and/or decreases in the spectral gap $\gamma$ (weaker separation between $\lambda_{1}$ and the rest). Intuitively, when a low-degree vertex is embedded into a hub-centered local structure, its influence is no longer well summarized by its degree alone: it can ‘borrow’ the hub’s amplification along walks, so perturbations supported on such mixed hub–periphery motifs are more easily transmitted across modes. The resulting modal interaction is exactly the regime in which $\kappa$ can grow and $\gamma$ can shrink [^16] In general $\kappa(A)$ and $\gamma(A)$ are not invariant under degree-preserving rewiring: they depend on the current adjacency matrix and may drift along the trajectory. Accordingly, as in earlier remarks, we interpret the bound uniformly on a reachable region $S_{0}$, working with time-invariant descriptors $\kappa^{\ast}:=\sup_{A\in S_{0}}\kappa(A)$ and $\gamma^{\ast}:=\inf_{A\in S_{0}}\gamma(A)>0$. In heavy-tailed settings these suprema/infima can be larger (worse), since hub-centered local-structure formation can increase modal interaction (raising $\kappa(A)$ and potentially shrinking $\gamma(A)$)..
These considerations point to two opposing forces. The direct effect of heavy tails is leverage: for a fixed perturbation budget they typically generate larger changes in the statistic—equivalently, for a fixed change in the statistic they can require fewer accepted swaps and hence permit smaller angle deviation. The indirect effect is spectral: if hub-centered structures worsen conditioning (larger $\kappa$ and/or smaller $\gamma$), then a given perturbation can translate into a larger eigenvector rotation—equivalently, a fixed change in the statistic can be accompanied by a larger angle deviation. The net effect of degree-tail heaviness on the angle–moment bounds is therefore a priori ambiguous, and depends on which channel dominates in the regime of interest.
@@ 7 Concluding remarks
Applied researchers working in data-scarce environments seldom have the information necessary to compute the eigenvectors of the network. Since the eigenvector is a good description of the systemic significance of nodes—whether it be firms about to foreclose or humans about to transmit a disease—the researcher is left with having to proxy it. Which he typically does with the degree vector [^17] Typically, the number of connections of a node is more easily known than the entire network structure.. Over the years, a certain ‘folk theorem’ has emerged around this substitution, which claims that the procedure is not all too bad when the network has neutral degree-mixing and lacks meso structures. This folk theorem, however, does not tell us anything about the error introduced by the substitution in networks that violate the neutrality assumptions. This paper presents analytical bounds on the error that is born from substituting the degree-vector for the eigenvector in assortative networks with a modicum of meso structures. We start from a neutral benchmark in which degree and eigenvector centrality align, and then introduce four common departures from neutrality: degree assortativity, community structure, core-periphery organization, and short directed cycles. For each departure, we describe a degree-preserving rewiring mechanism that strengthens the corresponding structure, construct the induced perturbation of the adjacency matrix, and translate it into an upper bound on the deviation angle between the degree vector and the leading eigenvector via the Stewart-Sun Perturbation Bound. With this procedure, we placed a boundary around an error term of some empirical importance in economics, epidemiology, opinion dynamics, and other applied sciences.

### v2

@@ 6 Power-law tails and statistic-specific bounds
For assortativity, the relevant constants are governed by endpoint-degree dispersion. Newman assortativity is a normalized covariance across edges, and $\phi_{p,q}$ correlates the $p$-degree at the tail of an edge with the $q$-degree at its head. A degree-preserving swap cannot alter the endpoint-degree multisets; it can only change which tails are paired with which heads. That is the only margin along which assortativity moves. The normalization
$\nu_{p,q}=\frac{1}{\sigma_{T}\sigma_{H}}$
rescales the raw degree-product covariance into a dimensionless correlation, where $\sigma_{T}$ and $\sigma_{H}$ are the standard deviations of the tail- and head-side endpoint-degree samples. Along the rewiring trajectory these dispersions, and hence $\nu_{p,q}$, remain fixed, so rewiring acts only through changes in cross-edge covariance. Because each swap re-pairs only a constant number of endpoints, the per-swap movement in $\phi_{p,q}$ is driven by the upper tail. Most swaps touch only moderate-degree vertices and have little effect, but a swap involving a rare high-degree node can generate a much larger jump in the degree-product sum. In the power-law regime one has
$\Lambda_{p,q}\asymp\nu_{p,q}d_{\max}^{2}$
and hence
$M_{p,q}\;\lesssim\;\widetilde{C}_{p,q}\,\nu_{p,q}\,\frac{\kappa_{*}}{\gamma_{*}}\,r\,n^{2/\alpha}.$
Thus smaller $\alpha$, corresponding to heavier tails, enlarges the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ by increasing the leverage of extremal endpoint degrees.
Communities
For community structure, the constants are shaped by the chosen partition because $\phi_{\mathrm{com}}$ records how much edge mass lies within blocks rather than across them. A degree-preserving swap cannot change how many stubs each vertex contributes, so the only way to increase $\phi_{\mathrm{com}}$ is to reroute existing cross-block mass into within-block locations while keeping all row and column sums fixed. How fast this can happen depends on two ingredients. First, the partition geometry matters. Smaller or more numerous communities provide less within-block capacity, so the process saturates earlier. Larger or more balanced blocks provide more admissible within-block destinations. Second, degree heterogeneity matters because most of the movable mass is carried by a relatively small set of high-degree vertices. If these hubs have many admissible within-block partners, cross-block mass can be converted quickly; if not, gains in $\phi_{\mathrm{com}}$ stall even when many swaps remain available. The affine constants therefore depend jointly on partition geometry and the degree tail: the partition determines the available within-block capacity, while the tail determines how concentrated the movable mass is. In the power-law regime,
$\Lambda_{\mathrm{com}}\asymp C_{\mathrm{com}}(\mathcal{C})\,d_{\max}^{2}$
and therefore
$M_{\mathrm{com}}\;\lesssim\;\widehat{C}_{\mathrm{com}}(\mathcal{C})\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\mathrm{com}}\,n^{2/\alpha}.$
Hence heavier tails enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ through greater extremal-degree leverage under degree-preserving rewiring.
Core–periphery
Core–periphery has a similar mass-transfer logic, but the target is asymmetric. The statistic $\phi_{\mathrm{cp}}$ rises primarily by draining mass from the $\mathcal{L}\!\to\!\mathcal{L}$ block and redirecting it so that periphery endpoints connect through the core $\mathcal{H}$. Since vertexwise in- and out-degrees are fixed, rewiring cannot change how much peripheral stub mass exists; it can only change where that mass lands. The relevant constants are therefore governed by the split $(\mathcal{H},\mathcal{L})$ and by the absorptive capacity of the core. If the core has enough degree mass, periphery–periphery edges can be progressively reattached through $\mathcal{H}$; if not, some $\mathcal{L}\!\to\!\mathcal{L}$ residue is unavoidable. Degree dispersion matters because in heavy-tailed networks a few very high-degree core vertices act as large-capacity hubs, allowing more peripheral mass to be redirected before saturation occurs. This gives core–periphery relatively high leverage: a modest number of well-placed swaps can rapidly concentrate paths through the core, to which the Perron vector is especially responsive. In the power-law regime,
$\Lambda_{\mathrm{cp}}\asymp C_{\mathrm{cp}}(\mathcal{H},\mathcal{L})\,d_{\max}^{2}$
and hence
$M_{\mathrm{cp}}\;\lesssim\;\widehat{C}_{\mathrm{cp}}(\mathcal{H},\mathcal{L})\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{\mathrm{cp}}\,n^{2/\alpha}.$
So here too heavier tails enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ through extremal-degree leverage under degree-preserving rewiring.
Cycles
@@
For $k$-cycles, $\phi_{k\text{-}\mathrm{cyc}}$ is motif-based, so the relevant rewiring is local. A single accepted swap can create a new short cycle, or complete an incipient one, by modifying only a constant number of edges. This locality makes cycles a useful stress test for eigenvector robustness: a small motif inserted in a small region may induce noticeable localized eigenvector distortion even when coarse global summaries move slowly. Degree heterogeneity again matters combinatorially. High-degree vertices provide many potential neighbors, so swaps involving hubs often make it easier to close cycles because there are more admissible ways to complete the required adjacency pattern under degree preservation. Longer cycles are usually costlier because building a $k$-cycle requires coordinating adjacency among $O(k)$ vertices and often multiple accepted swaps. But heavy tails mitigate this cost by supplying hubs as attachment points, making long-cycle assembly far easier than in thin-tailed graphs. Accordingly, in the power-law regime,
$\Lambda_{k\text{-}\mathrm{cyc}}\asymp C_{k}d_{\max}^{2}$
and therefore
$M_{k\text{-}\mathrm{cyc}}\;\lesssim\;\widehat{C}_{k}\,\frac{\kappa_{*}}{\gamma_{*}}\,r_{k\text{-}\mathrm{cyc}}\,n^{2/\alpha}.$
Thus heavier tails enlarge the time-uniform envelope for $\mathbb{E}[\theta_{t}]$ by increasing the combinatorial leverage available for closing cycles around hubs.
Across all four cases, the same mechanism is at work. Heavy degree tails determine how much structural change can be generated with a fixed degree-preserving perturbation budget because a small set of hubs carries a disproportionate share of edge-ends. This makes the evolution uneven: swaps involving hubs can move the statistic substantially, whereas most swaps that avoid them do relatively little. The displayed bounds capture this through the common scaling
$M\;\lesssim\;(\text{statistic-specific constant})\cdot\frac{\kappa_{*}}{\gamma_{*}}\cdot r\cdot n^{2/\alpha},$
where the statistic-specific constant reflects whether one is re-pairing endpoints, redirecting mass across blocks, draining periphery–periphery edges through the core, or closing local motifs. The factor $n^{2/\alpha}$ is the signature of extremal-degree leverage.
There is, however, a second channel through which heavy tails can matter. They may also affect the spectral sensitivity parameters themselves. Hub-centered rewiring tends to build several forms of structure around the same small set of vertices, which can worsen eigenbasis conditioning or reduce modal separation. In the language of the Stewart–Sun bound, this means that heavy-tailed rewiring may increase $\kappa$ and/or reduce $\gamma$, amplifying the eigenvector response to a given perturbation. [^16] In general $\kappa(A)$ and $\gamma(A)$ are not invariant under degree-preserving rewiring: they depend on the current adjacency matrix and may drift along the trajectory. Accordingly, as in earlier remarks, we interpret the bound uniformly on a reachable region $S_{0}$, working with the time-invariant descriptors $\kappa_{*}:=\sup_{A\in S_{0}}\kappa(A),\qquad\gamma_{*}:=\inf_{A\in S_{0}}\gamma(A)>0.$ In heavy-tailed settings these worst-case descriptors can deteriorate, since hub-centered local structure can increase modal interaction and weaken eigenvalue separation. Heavy tails therefore operate through two distinct channels. The direct channel is combinatorial: they increase the leverage with which degree-preserving swaps can move the chosen statistic. The indirect channel is spectral: they may also make the leading eigenvector more fragile by worsening conditioning or shrinking separation. The net effect on the angle–moment bounds is therefore not mechanically monotone; it depends on which channel dominates in the regime under study.
@@ 7 Concluding remarks
Applied researchers working in data-scarce environments seldom have the information necessary to compute the eigenvectors of the network. Since the eigenvector is a good description of the systemic significance of nodes—whether it be firms about to foreclose or humans about to transmit a disease—the researcher is left with having to proxy it. Which he typically does with the degree vector [^17] Typically, the number of connections of a node is more easily known than the entire network structure.. Over the years, a certain ‘folk theorem’ has emerged around this substitution, which claims that the procedure is not all too bad when the network has neutral degree-mixing and lacks meso structures. This folk theorem, however, does not tell us anything about the error introduced by the substitution in networks that violate the neutrality assumptions. This paper presents analytical bounds on the error that is born from substituting the degree-vector for the eigenvector in assortative networks with a modicum of meso structures. We start from a neutral benchmark in which degree and eigenvector centrality align, and then introduce four common departures from neutrality: degree assortativity, community structure, core-periphery organization, and short directed cycles. For each departure, we describe a degree-preserving rewiring mechanism that strengthens the corresponding structure, construct the induced perturbation of the adjacency matrix, and translate it into an upper bound on the deviation angle between the degree vector and the leading eigenvector via the Stewart-Sun Perturbation Theorem. We broaden this methodology to encompass weighted networks as well, showing that continuous strength-preserving weight transfers yield analogous analytical bounds. Furthermore, our theoretical framework is complemented by numerical simulations that illustrate the real-time dynamics of eigenvector rotation, explicitly mapping the wandering and targeted drifting of the deviation angle under various structural perturbations. With this procedure, we placed a boundary around an error term of some empirical importance in economics, epidemiology, opinion dynamics, and other applied sciences.